OpenStudy (lgbasallote):

a thermometer reading 18 F is brought into a room, the temperature of which is 70 F. one minute later the reading is 31 F. determine the temperature reading as a function of time. find the temperatire reading 5 minutes after the thermometer is first brought into the room

OpenStudy (lgbasallote):

Given T_m = 70 F @ t = 0 T = 18 @ t = 1 T = 31

OpenStudy (lgbasallote):

\[\ln (\frac{31 - 70}{18-70}) = k\] \[\ln (0.75) = k\] \[k = -0.2877\]

OpenStudy (lgbasallote):

when t = 5 then \[\ln (\frac{T - 70}{18 - 70} )= -0.2877(5)\] \[\ln (\frac{T-70}{-52}) = -1.4385\] \[T = 70 - 52e^{-1.4385}\] \[T = 57.66 F\]

OpenStudy (lgbasallote):

is that right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

T as a function of time \[\large T=70-52 e^{-1.4t}\]

OpenStudy (lgbasallote):

okay thanks :D